# Med-GLIP: Advancing Medical Language-Image Pre-training with Large-scale Grounded Dataset

### v1

@@ IV-B Results & Analysis
After full finetuning, Med-GLIP outperforms CO-DETR in five out of seven modalities, including CT, MRI, X-ray, ultrasound, and dermoscopy. For instance, Med-GLIP achieves substantially higher AP scores on CT (82.7 vs. 31.8), MRI (28.0 vs. 7.6), X-ray (86.9 vs. 69.8), ultrasound (59.2 vs. 1.0), and dermoscopy (61.7 vs. 43.0). Although CO-DETR attains slightly better performance in endoscopy and fundus photography, Med-GLIP demonstrates robust and stable performance across a wider range of modalities, highlighting its strong generalizability and effectiveness as a unified framework for medical grounding tasks. Overall, these results demonstrate that our dataset and adaptation strategy benefit a variety of grounding models, while Med-GLIP in particular achieves the best performance (Figure 5).

### v2

@@ IV-B Results & Analysis
After full finetuning, Med-GLIP outperforms CO-DETR in five out of seven modalities, including CT, MRI, X-ray, ultrasound, and dermoscopy. For instance, Med-GLIP achieves substantially higher AP scores on CT (82.7 vs. 31.8), MRI (28.0 vs. 7.6), X-ray (86.9 vs. 69.8), ultrasound (59.2 vs. 1.0), and dermoscopy (61.7 vs. 43.0). Although CO-DETR attains slightly better performance in endoscopy and fundus photography, Med-GLIP demonstrates robust and stable performance across a wider range of modalities, highlighting its strong generalizability and effectiveness as a unified framework for medical grounding tasks. Overall, these results demonstrate that our dataset and adaptation strategy benefit a variety of grounding models, while Med-GLIP in particular achieves the best performance (Table II).
@@ Conclusion
Acknowledgment
This work is supported by the National Key R&D Program of China (Grant No. 2024YFC3308304), the ”Pioneer” and ”Leading Goose” R&D Program of Zhejiang (Grant No. 2025C01128), the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008).